# Extragradient methods with complexity guarantees for hierarchical variational inequalities

### v1

@@ 5.2 Tseng splitting
Optimistic extragradient (also known as Popov’s or past extragradient) method saves one evaluation of the time-varying operator $V_{k}$ per iteration compared to the standard extragradient method. Yet, it still requires two evaluations of the proximal operator of $G_{k}$ per iteration. An attractive modified extragradient scheme, which uses past values of the operator but saves on one evaluation of the proximal operator, is the following optimistic version of Tseng’s Forward-Backward-Forward method [Tsen00, tran2024revisiting]
@@
$z^{k+1/2}$ $=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k}),$
$z^{k+1}$ $=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k})-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}))$
$=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k}).$
This further gives us
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k+1}-z^{k}+z^{k}-z}^{2}$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+\langle z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k})-z^{k},z^{k+1}-z\rangle$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
$\hskip 17.00024pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle.$ (5.21)
@@
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})$ $+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
$=2\langle z^{k}-z^{k+1/2},z^{k+1}-z^{k+1/2}\rangle+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle$
$=\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1}-z^{k}}$
$\hskip 17.00024pt2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle.$
@@
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k}}^{2}$
$\hskip 17.00024pt2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle$
$\hskip 17.00024pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi_{k},z^{k+1/2}-z\rangle$
Since $\xi^{k}\in\partial G_{k}(z^{k+1/2})$ it holds that $G_{k}(z^{k+1/2})-G_{k}(z)\leq\langle\xi^{k},z^{k+1/2}-z\rangle$ and therefore
$-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle$ $=-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle-2t_{k}\langle\xi^{k},z^{k+1/2}-z\rangle$
$\leq 2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+2t_{k}(G_{k}(z)-G_{k}(z^{k+1/2}))$ (5.22)
Moreover, Young’s inequality yields
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle\leq t_{k}^{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\norm{z^{k+1}-z^{k+1/2}}.$
With this estimate, we arrive at (4.1)
$\frac{1}{2}\norm{z^{k+1}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ (5.23)
$\hskip 17.00024pt+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}(G(z)-G(z^{k+1/2})).$
Repeating the same steps as in the rest of Section 4.1, we obtain the same energy bound (4.12). Using this energy bound and repeating the same arguments as in Section 4, we obtain the same results as in Theorem 3.1 and Corollary 3.2.
@@ 6.1 Hierarchical Nash equilibria
To provide a simple illustration on the practical performance of our method, we’ve tested Algorithm 1 on a simple version of the hierarchical equilibrium problem described Section 1.1.4, taken from [LamSagSIOPT25]. We’ve chosen a deliberately simple example for which the equilibrium set of the game can be computed explicitly and we can monitor the evolution of the inexactness of the algorithm. Let us consider $N=4$ lower-level players and $M=2$ upper-level players, with $x^{1}=(y^{2},y^{4}),x^{2}=(y^{1},y^{3})$. The player-cost functions are
$h^{\ell}_{1}(y^{1},y^{-1})=0.5(y^{1})^{2}+y^{1}(y^{2}+2y^{3}+y^{4}-100),\;\varphi^{\ell}_{1}(y^{1})=\delta_{[-100,50]}(y^{1}),$
$h^{\ell}_{2}(y^{2},y^{-2})=0.5(y^{2})^{2}+y^{2}(y^{1}+y^{3}+y^{4}-50),\;\varphi^{\ell}_{2}(y^{2})=\max\{-10(y^{2}-15),0\}+\delta_{[0,50]}(y^{2}),$
$h^{\ell}_{3}(y^{3},y^{-3})=0.5(y^{3})^{2}+y^{3}(y^{2}+y^{4}-100),\;\varphi^{\ell}_{3}(y^{3})=\delta_{[0,100]}(y^{3})$
$h^{\ell}_{4}(y^{4},y^{-4})=0.5(y^{4})^{2}+y^{4}(y^{1}+y^{2}+y^{2}+y^{3}-50),\;\varphi^{\ell}_{4}(y^{4})=\delta_{[0,50]}(y^{4}).$
@@
$h^{u}_{2}(x_{1},x_{2})=(y^{1})^{2}+y^{1}(y^{2}+y^{3})+(y^{3})^{2}+y^{3}(y^{2}+y^{4}),\varphi^{u}_{2}(x^{1})=0.$
One can obtain an explicit expression for the lower-level equilibrium set: $\mathcal{S}_{2}=\{(-50,y^{2},50,50-y^{2}):15\leq y^{2}\leq 50\}$. Thus, the unique variational equilibrium of this game is (50,15,50,35)(-50,15,50,35). We’ve solved this game problem with Algorithm 1 using the regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size. Figure 1 displays the temporal evolution of the strategies of the four players, as well as the distance to the unique variational equilibrium
Figure 1: Evolution of the strategy profile and error plot.
@@ 6.2 Constrained Min-Max problem
where ${\mathbf{A}}:\mathcal{X}\to\mathcal{Z}?$ and ${\mathbf{B}}:\mathcal{Y}\to\mathcal{Z}?$ are bounded linear operators. The saddle function $\mathcal{L}:\mathcal{Z}\triangleq\mathcal{X}\times\mathcal{Y}\to(-\infty,\infty]$ is of the form
@@
Using the product space structure $z\triangleq(x,y)\in\mathcal{X}\times\mathcal{Y}\triangleq\mathcal{Z}$, the optimistic extragradient method can be applied to solve this family of equilibrium problems.
Generalized Absolute value equations (GAVEs).
The Generalized Absolute value equations (GAVEs) [mangasarian2006absolute] is an important non-smooth NP-hard problem in the form of
@@
For our numerical investigations, we’ve implemented the following example from [Ebadi24]: Consider the boundary value problem
@@
$u(0)=-1,u(1)=0$
@@
Figure 2: Evolution of the approximate solution $u(k)$ after completion of the $k$-th iteration of the Algorithm (left), and the evolution of the residual relative to the true solution (right).
@@
We’ve reconstructed the solution $u(x)$ using Algorithm 1 with regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size. Figure LABEL:fig:GAVEODESolution displays the evolution of approximate solutions obtained with our method measured at the indicated snapshots of the algorithm.
@@ 7 Conclusion
This paper proves seminal results on the iteration complexity of hierarchical hemi-variational inequalities. Constructing a suitably defined optimistic extragradient method, we generalize the state-of-the art along several dimensions. In particular, we derive complexity guarantees in terms of suitably defined feasibility and optimality gaps. Our analysis reveals close connections to geometric conditions imposed on the operator defining the lower level equilibrium problems, conditions which have been so-far dominantly been used in the potential case. We belief that our proof strategy is going to be useful for several other problems currently under investigation in the literature. In particular, it will be very interesting to derive rates for stochastic formulation of our model template and derive rates in terms of expected gap functions. Further interesting directions for future research inlude the consideration of accelerated methods. We leave these important extensions to future research.
@@ Acknowledgements
This research benefited from the support of the FMJH Program PGMO. MST’s research is supported by the Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 556222748 "non-stationary hierarchical minimization". We thank Johannes-Carl Schnebel for his help on conducting the numerical experiments, and Enis Chenchene, Radu I. Bot, as well as David A. Hulett for fruitful discussions on this topic.
@@
Appendix A The Fitzpatrick function
Let $\mathsf{M}:\mathcal{Z}\to 2^{\mathcal{Z}}$ be a maximally monotone operator. The Fitzpatrick function [fitzpatrick1988representing, BauCom16] $\mathcal{F}_{\mathsf{M}}:\mathcal{Z}\to(-\infty,+\infty]$, associated with the operator $\mathsf{M}$, is defined as
@@ Appendix A The Fitzpatrick function
Following [Borwein:2016aa], we define the gap function $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\triangleq\mathcal{F}_{\mathsf{M}}(x,0)$. $\operatorname{\mathsf{Gap}}_{\mathsf{M}}$ is convex, and in fact the smallest translation invariant gap function associated with the monotone operator $\mathsf{M}$ [Borwein:2016aa, Theorem 3.1]. Importantly, this gives the properties $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\geq 0$ and $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)=0$ if and only if $x\in\operatorname{zer}(\mathsf{M})$. To make this concept concrete, observe that if $\mathsf{M}={\mathsf{F}}+\operatorname{\mathsf{NC}}_{\mathcal{C}}$, then the above definition of the gap function reduces to the well-known Auslender dual gap function [FacPan03, auslender1976optimisation]
@@
To the data $({\mathsf{F}},g)$, we thus associate the bifunction $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ defined as
$H^{({\mathsf{F}},g)}(x,y)\triangleq\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y).$
If ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is monotone and continuous on $\operatorname{dom}(g)$, it is easy to verify that $H^{({\mathsf{F}},g)}(x,y)\leq-H^{({\mathsf{F}},g)}(y,x)$ for all $(x,y)\in\operatorname{dom}(g)\times\operatorname{dom}(g)$ (i.e. $H^{({\mathsf{F}},g)}$ is a monotone bifunction [iusem2011maximal]). In the structured setting $\operatorname{HVI}({\mathsf{F}},g)$, we obtain the following bounds on the Fitzpatrick function, showing its close connection to the bifunction $H^{({\mathsf{F}},g)}$ and the Auslender dual gap function. First, the convex subgradient inequality yields the relation
@@
$\varphi^{({\mathsf{F}},g)}(x,u)\leq\sup_{y\in\operatorname{dom}(g)}\{\langle y,u\rangle-H^{({\mathsf{F}},g)}(y,x)\}=\left(H^{({\mathsf{F}},g)}(\bullet,x)\right)^{\ast}(u).$ (A.1)
@@
$\varphi^{({\mathsf{F}},g)}(x,u)\geq\langle x,u\rangle\hskip 17.00024pt\forall x\in\operatorname{dom}(g),$ (A.2)
$\mathcal{F}_{{\mathsf{F}}+\partial g}(x,0)\leq\sup_{y\in\operatorname{dom}(g)}H^{({\mathsf{F}},g)}(x,y)=\varphi^{({\mathsf{F}},g)}(x,0).$ (A.3)
Following this notation, we define for a given a subset $\mathcal{C}\subseteq\mathcal{Z}$, the restricted dual gap function as
$\Theta(x|{\mathsf{F}},g,\mathcal{C})\triangleq\sup_{y\in\mathcal{C}}H^{({\mathsf{F}},g)}(x,y).$ (A.4)
Appendix B Omitted proofs

### v2

@@ 5.2 Tseng splitting
Optimistic extragradient (also known as Popov’s or past extragradient) method saves one evaluation of the time-varying operator ${\mathsf{V}}_{k}$ per iteration compared to the standard extragradient method. Yet, it still requires two evaluations of the proximal operator of $G_{k}$ per iteration. An attractive modified extragradient scheme, which uses past values of the operator but saves on one evaluation of the proximal operator, is the following optimistic version of Tseng’s Forward-Backward-Forward method [62, 61]
@@
$z^{k+1/2}=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k}),\text{ and }z^{k+1}=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k}).$
Eq. (4.8) gives
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle$
$\hskip-42.50061pt=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+\langle z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k})-z^{k},z^{k+1}-z\rangle$
$\hskip-42.50061pt=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
$\hskip-25.50037pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle.$ (5.15)
@@
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1}-z^{k+1/2}\rangle=2\langle z^{k}-z^{k+1/2},z^{k+1}-z^{k+1/2}\rangle$
$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle$
$=\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1}-z^{k}}^{2}$
$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle.$
@@
$\norm{z^{k+1}-z}^{2}=\norm{z^{k}-z}^{2}-\norm{z^{k}-z^{k+1/2}}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}$
$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi_{k},z^{k+1/2}-z\rangle.$
Since $\xi^{k}\in\partial G_{k}(z^{k+1/2})$ the subgradient inequality allows us to continue with
$\norm{z^{k+1}-z}^{2}$ $\leq\norm{z^{k}-z}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}$
$-2t_{k}\Psi_{k}(z)-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle.$
where we have used the notation introduced in (4.2). This is exactly (4.10). Repeating the same steps as in the rest of Section 4.1, we obtain the same energy bound (4.5). Using this energy bound and repeating the same arguments as in Section 4, we obtain the same results as reported in Section 3.1 for the optimistic extragradient method.
@@ 6.1 Hierarchical Nash equilibria
To provide a simple illustration on the practical performance of our method, we’ve tested Algorithm 1 on a simple version of the hierarchical equilibrium problem described in Section 1.1.3, taken from [38]. We’ve chosen a deliberately simple example for which the equilibrium set of the game can be computed explicitly and we can monitor the evolution of the inexactness of the algorithm. Let us consider $N=4$ lower-level players and $M=2$ upper-level players, with $x^{1}=(y^{2},y^{4}),x^{2}=(y^{1},y^{3})$. The player-cost functions are
$h^{\ell}_{1}(y^{1},y^{-1})=0.5(y^{1})^{2}+y^{1}(y^{2}+2y^{3}+y^{4}-100),\;\varphi^{\ell}_{1}(y^{1})=\iota_{[-100,50]}(y^{1}),$
$h^{\ell}_{2}(y^{2},y^{-2})=0.5(y^{2})^{2}+y^{2}(y^{1}+y^{3}+y^{4}-50),$
$\varphi^{\ell}_{2}(y^{2})=\max\{-10(y^{2}-15),0\}+\iota_{[0,50]}(y^{2}),$
$h^{\ell}_{3}(y^{3},y^{-3})=0.5(y^{3})^{2}+y^{3}(y^{2}+y^{4}-100),\;\varphi^{\ell}_{3}(y^{3})=\iota_{[0,100]}(y^{3})$
$h^{\ell}_{4}(y^{4},y^{-4})=0.5(y^{4})^{2}+y^{4}(y^{1}+y^{2}+y^{2}+y^{3}-50),\;\varphi^{\ell}_{4}(y^{4})=\iota_{[0,50]}(y^{4}).$
@@
$h^{u}_{2}(x_{1},x_{2})=(y^{1})^{2}+y^{1}(y^{2}+y^{3})+(y^{3})^{2}+y^{3}(y^{2}+y^{4}),\varphi^{u}_{2}(x^{2})=0.$
One can obtain an explicit expression for the lower-level equilibrium set: $\mathcal{S}_{2}=\{(-50,y^{2},50,50-y^{2}):15\leq y^{2}\leq 50\}$. Thus, the unique variational equilibrium of this game is (50,15,50,35)(-50,15,50,35). We have solved this game problem with Algorithm 1 using the regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size $t_{k}=t$. Figure 1 displays the temporal evolution of the strategies of the four players, as well as the distance to the unique variational equilibrium.
Figure 1: Evolution of the strategy profile and error plot for the hierarchical NEP.
@@ 6.2 Constrained Min-Max problem
where ${\mathbf{A}}:\mathcal{X}\to\mathcal{Z}$ and ${\mathbf{B}}:\mathcal{Y}\to\mathcal{Z}$ are bounded linear operators. The saddle function $\mathcal{L}:\mathcal{Z}\triangleq\mathcal{X}\times\mathcal{Y}\to(-\infty,\infty]$ is of the form
@@
Using the product space structure $z\triangleq(x,y)\in\mathcal{Z}$, the optimistic extragradient method can be applied to solve this family of equilibrium problems.
Generalized Absolute value equations (GAVEs)
The GAVEs [43] is an important non-smooth NP-hard problem in the form of
@@
For our numerical investigations, we’ve implemented the following example from [23]. Consider the boundary value problem
[…]
$u(0)=-1,u(1)=0,$
@@
Figure 2: Evolution of the approximate solution $u^{k}$ after completion of the $k$-th iteration of the Algorithm (left), and the evolution of the residual relative to the true solution (right).
@@
We’ve reconstructed the solution $u(x)$ using Algorithm 1 with regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size $t_{k}=t$. Figure 2 displays the evolution of approximate solutions obtained with our method measured at the indicated snapshots of the algorithm.
@@ 7 Conclusion
This paper proves several results on the convergence rates for solving hierarchical HVI’s. Constructing a suitably defined optimistic extragradient method, we advance the state of the art along several dimensions. In particular, we derive convergence rates in terms of suitably defined feasibility and optimality gaps. Our analysis reveals close connections to geometric conditions imposed on the operator defining the lower level equilibrium problems, conditions which have so-far only been used in the context of hierarchical optimization. We believe that our proof strategy is going to be useful for several other problems currently under investigation in the literature. In particular, it will be very interesting to derive rates for the stochastic formulation of our model template and derive rates in terms of expected gap functions. Further interesting directions for future research include the consideration of accelerated methods. We leave these questions as important extensions to future research.
@@ Acknowledgements
This research benefited from the support of the FMJH Program PGMO. MST’s research is supported by the Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 556222748 "non-stationary hierarchical minimization". PD’s research is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689).
We thank Johannes-Carl Schnebel for his help on conducting the numerical experiments, and Enis Chenchene, Radu I. Bot, as well as David A. Hulett for fruitful discussions on this topic.
@@
Appendix A Auxiliary facts
Lemma A.1 (Lemma 5.31, [9]).
Let $(a_{k})_{k\geq 1},(b_{k})_{k\geq 1},(c_{k})_{k\geq 1}$ be nonnegative sequences such that $\sum_{k\geq 1}c_{k}<\infty$ and $a_{k+1}\leq a_{k}-b_{k}+c_{k}.$ Then, $\lim_{k\to\infty}a_{k}$ exists and $\sum_{k\geq 1}b_{k}<\infty$.
Let $(x_{n})_{n\geq 1}$ be a sequence in $\mathcal{Z}$ and $(\lambda_{n})_{n\geq 1}$ a sequence of positive numbers such that $\sum_{n\geq 1}\lambda_{n}=+\infty$. Define the sequence of weighted averages $z_{n}=\frac{1}{\tau_{n}}\sum_{k=1}^{n}\lambda_{k}x_{k}$, where $\tau_{n}\triangleq\sum_{k=1}^{n}\lambda_{k}$. The next fact is Lemma 2.3 in [6].
Lemma A.2 (Opial-Pasty).
Let $\mathcal{S}$ be a nonempty subset of $\mathcal{Z}$ and assume that $\lim_{n\to\infty}\norm{x_{n}-x}$ exists for every $x\in\mathcal{S}$. If every weak cluster point of $(x_{n})_{n\geq 1}$ (respectively $(z_{n})_{n\geq 1}$) lies in $\mathcal{S}$, then $(x_{n})_{n\geq 1}$ (respectively $(z_{n})_{n\geq 1}$) converges weakly to an element of $\mathcal{S}$ as $n\to\infty$.
Appendix B The Fitzpatrick function
Let $\mathsf{M}:\mathcal{Z}\to 2^{\mathcal{Z}}$ be a maximally monotone operator. The Fitzpatrick function [26, 9] $\mathcal{F}_{\mathsf{M}}:\mathcal{Z}\to(-\infty,+\infty]$, associated with the operator $\mathsf{M}$, is defined as
@@ Appendix B The Fitzpatrick function
Following [12], we define the gap function $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\triangleq\mathcal{F}_{\mathsf{M}}(x,0)$. $\operatorname{\mathsf{Gap}}_{\mathsf{M}}$ is convex, and in fact the smallest translation invariant gap function associated with the monotone operator $\mathsf{M}$ [12, Theorem 3.1]. Importantly, this gives the properties $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\geq 0$ and $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)=0$ if and only if $x\in\operatorname{zer}(\mathsf{M})$. To make this concept concrete, observe that if $\mathsf{M}={\mathsf{F}}+\operatorname{\mathsf{NC}}_{\mathcal{C}}$, then the above definition of the gap function reduces to the well-known Auslender dual gap function [25, 7]
@@
Recall the definition (2.5) of the bifunction $H^{({\mathsf{F}},g)}$. If ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is monotone and continuous on $\operatorname{dom}(g)$, it is easy to verify that $H^{({\mathsf{F}},g)}(x,y)\leq-H^{({\mathsf{F}},g)}(y,x)$ for all $(x,y)\in\operatorname{dom}(g)\times\operatorname{dom}(g)$ (i.e. $H^{({\mathsf{F}},g)}$ is a monotone bifunction [32]). In the structured setting $\operatorname{HVI}({\mathsf{F}},g)$, we obtain the following bounds on the Fitzpatrick function, showing its close connection to the bifunction $H^{({\mathsf{F}},g)}$ and the Auslender dual gap function. First, the convex subgradient inequality yields the relation
@@
$\varphi^{({\mathsf{F}},g)}(x,u)\leq\sup_{y\in\operatorname{dom}(g)}\{\langle y,u\rangle-H^{({\mathsf{F}},g)}(y,x)\}=\left(H^{({\mathsf{F}},g)}(\bullet,x)\right)^{\ast}(u).$ (B.1)
@@
$\varphi^{({\mathsf{F}},g)}(x,u)\geq\langle x,u\rangle\hskip 17.00024pt\forall x\in\operatorname{dom}(g),$ (B.2)
$\mathcal{F}_{{\mathsf{F}}+\partial g}(x,0)\leq\sup_{y\in\operatorname{dom}(g)}H^{({\mathsf{F}},g)}(x,y)=\varphi^{({\mathsf{F}},g)}(x,0).$ (B.3)
Following this notation, we define the restricted dual gap function as in (2.4).
Appendix C Omitted proofs